3. Calculating Taxes. The Drydel Co. had $307,000 in taxable income. Using the tax rates from

the chart provided. Calculate the company’s income taxes. Show your computation.Tax Rate 15% Taxable income 0- 50,000 50,001 - 75,000 75,001 - 100,000 100,001 - 335,000 335,001 -10,000,000 10,000,001 - 15,
Answer #1

Taxes will be as follows

On first 50,000 =50,000*15% =$7,500

Next 25,000 =25,000*25% =$6,250

Next 25,000 =25000*34% =$8,500

Next 207,000 =207,000*39% =$80,730

Hence, total income tax =$102,980